Find the missing dimension in each rectangle. Length is 13m and perimeter is 68m
Answer by checkley77(12844)

2*13+2*W=68
26+2W=68
2W=68-26
2W=42
W=42/2
W=21M.
PROOF:
2+13+2*21=68
26+42=68
68=68